# Curvature of basis exchange walks

### v1

@@ 1 Introduction
Recent works have shown the basis exchange walk on matroids have several nice properties: fast mixing [ALG+21], 1-spectral independence, and a lower bounded modified log Sobolev constant [CGM19]. These works take advantage of a crucial shared property for all basis exchange walks: a stationary distribution with a strongly log-concave generating polynomial (see [AOGV18] for more details). In this note, we identify one property that cannot be derived from this polynomial: non-negative Ollivier-Ricci curvature.
Ollivier–Ricci curvature on metric spaces was first introduced in [Oll09], as an analog to Ricci curvature on manifolds. This definition allowed for several natural extensions of results in geometry, such as Bonnet-Myers theorem and bounds on Laplacian eigenvalues with respect to curvature, both of which apply only to spaces with non-negative curvature. A key feature of curvature in geometry is the local-to-global phenomenon, where understanding the curvature, a strictly local notion, can be used to prove global features. Ollivier–Ricci curvature retains this feature in some cases. For example, non-negative Ollivier-Ricci curvature can be used to show bounds on mixing (see [LP17] on path coupling for more details). Considering this phenomenon and the work of [ALG+21] showing fast mixing for basis exchange walks, it is reasonable to ask if non-negative curvature could be an underlying feature of these walks.
This work shows this is not the case: despite the uniform fast-mixing results of [ALG+21] for all bases exchange walks, there exist both negatively and non-negatively curved basis exchange walks. Furthermore, we show even among standard classes of matroids, such as linear matroids or graphic matroids, there exist both matroids with negatively and non-negatively curved basis exchange walks. This work reinforces the idea that non-negative curvature is a somewhat delicate phenomenon, which cannot be guaranteed by a strongly log-concave stationary distribution or representability of the matroid.
Concurrent works [BCC+22] and [Liu21] connect non-negative Ollivier–Ricci curvature and spectral independence, two different ways of locally proving expansion. In particular, [Liu21] shows
“the existence of a ‘good’ coupling for any sufficiently ‘local’ dynamics with stationary distribution $\mu$ implies spectral independence for $\mu$,”
where a good coupling can be taken to be non-negative Ollivier-Ricci curvature. This work provides a counterexample to the converse: while all basis exchange walks satisfy 1-spectral independence, there exist examples that are negatively curved. In other words, there exist chains with one local certificate of expansion (spectral independence) but not another (non-negative Ollivier-Ricci curvature).
@@ 5 Upper bound on curvature
We now give an upper bound on the curvature of the basis exchange walk. In order to state this bound, we fix any $S\sim T$ and define the following notation.

### v2

@@ 1 Introduction
In this note, we consider the discrete Ollivier-Ricci curvature of a specific class of Markov chains, basis exchange walks on matroids. Ollivier–Ricci curvature on metric spaces was first introduced in [OLL09], as an analog to Ricci curvature on manifolds. This definition allowed for several natural extensions of results in geometry, such as Bonnet-Myers theorem and a lower bound on the spectral gap (both of which apply only to spaces with non-negative curvature). Notably, non-negative discrete curvature also implies bounds on the mixing time of a Markov chain (see [LP17] on path coupling for more details).
Basis exchange walks on matroids are a well-studied class of Markov chains which exhibit many desirable properties: lower bounded spectral gap and lower bounded modified log Sobolev constant [CGM19], spectral independence, and fast mixing [ALG+21]. We note that all of these properties are implied by non-negative Ollivier-Ricci curvature and pose the natural follow-up question:
Do all basis exchange walks on matroids have non-negative Ollivier-Ricci curvature?
We answer this question negatively: despite the many desirable properties uniformly exhibited by basis exchange walks, there exist both negatively and non-negatively curved basis exchange walks. Furthermore, we show even among standard classes of matroids, such as linear matroids or graphic matroids, there exist both matroids with negatively and non-negatively curved basis exchange walks. This work reinforces the idea that non-negative curvature is a somewhat delicate phenomenon, which cannot be guaranteed by a strongly log-concave stationary distribution (a feature of all basis exchange walks) or the representability of the matroid.
We highlight in particular the implication this has for the connection between spectral independence and Ollivier-Ricci curvature, both local certificates of expansion. Concurrent works [BCC+22] and [LIU21] prove non-negative Ollivier-Ricci curvature implies spectral independence (see either work for precise, quantified statements). The validity of the converse to this statement was not previously known, but this work disproves it: basis exchange walks are uniformly 1-spectrally independent, but we establish several examples with negative Ollivier-Ricci curvature.
@@ 5 Upper bound on curvature
We now give an upper bound on the curvature of the basis exchange walk (we note the main function of this bound will be to prove specific examples have negative curvature). In order to state this bound, we fix any $S\sim T$ and define the following notation.